40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 1 08:50:48 2022, yuta, Summary, LSC, FPMI with REFL/AS55 trial 
    Reply  Thu Jul 14 00:10:08 2022, yuta, Summary, LSC, FPMI with REFL/AS55 trial continued AS55_I.pngAS55_Q.pngREFL55_I.pngREFL55_Q.png
       Reply  Fri Jul 15 19:13:22 2022, Paco, Summary, LSC, FPMI with REFL/AS55 demod phase adjust 
          Reply  Fri Jul 15 22:36:04 2022, rana, Summary, LSC, FPMI with REFL/AS55 demod phase adjust 
          Reply  Mon Jul 18 16:39:07 2022, Paco, Summary, LSC, FPMI locking procedure using REFL55 and AS55 DARM_07_18_2022_FMPI.pdfCARM_07_18_2022_FPMI.pdfMICH_07_18_2022_FPMI.pdffpmi_darm_nb_2022_07.pdf
             Reply  Mon Jul 18 21:41:42 2022, Anchal, Summary, LSC, FPMI locking procedure using REFL55 and AS55 
             Reply  Tue Aug 9 19:54:31 2022, yuta, Summary, LSC, FPMI locking tonight 
                Reply  Thu Aug 18 14:49:35 2022, Yehonathan, Summary, LSC, FPMI Sensitivity Sensitivity_Plot_1344133503.pdf
                   Reply  Thu Aug 18 18:10:49 2022, Koji, Summary, LSC, FPMI Sensitivity 
Message ID: 17069     Entry time: Tue Aug 9 19:54:31 2022     In reply to: 17012     Reply to this: 17089
Author: yuta 
Type: Summary 
Category: LSC 
Subject: FPMI locking tonight 

[Tega, Anchal, Yuta]

We resored FPMI locking settings. Below is the summary of locking configurations tonight.
To ease the lock acquisition, the step to feedback POX11_I to ETMX and POY11_I to MC2 before POX and POY mixing was necessary tonight.

CARM (YARM):
 - 0.5 * POX11_I + 0.5 * POY11_I handed to 0.5 * REFL55_I
 - YARM filter module, FM4,5 for acquisition, FM1,2,3,6,8 triggered, C1:LSC-YARM_GAIN = 0.012
 - Actuation on -0.77 * MC2
 - UGF ~ 250 Hz

DARM (XARM):
 - 0.5 * POX11_I - 0.5 * POY11_I handed to 4.6 * AS55_Q (it was 2.5 in 40m/17012)
 - XARM filter module, FM5 for acquisition (no FM4), FM1,2,3,6,8 triggered, C1:LSC-XARM_GAIN = 0.015
 - Actuation on 0.5 * ETMX - 0.5 * ETMY
 - UGF ~ 120 Hz

MICH:
 - 1 * REFL55_Q (turned on after XARM and YARM acquisition)
 - MICH filter module, FM4,5,8 for acquisition, FM2,3 triggered, C1:LSC-MICH_GAIN = +40
 - Actuation on 0.5 * BS
 - UGF ~ 100 Hz

Measured sensing matrix:
Sensing Matrix with the following demodulation phases
{'AS55': 200.41785156862835, 'REFL55': 93.7514468401475, 'POX11': 105.08325063571438, 'POY11': -11.343909976281823}
Sensors              DARM                    CARM                   MICH
C1:LSC-AS55_I_ERR_DQ 5.27e-02 (-154.105 deg) 2.83e-01 (132.395 deg) 1.17e-04 (-40.1051 deg)
C1:LSC-AS55_Q_ERR_DQ 3.99e-02 (-151.048 deg) 1.42e-02 (125.504 deg) 1.41e-04 (-2.42846 deg)
C1:LSC-REFL55_I_ERR_DQ 5.59e-02 (77.6871 deg) 1.15e+00 (-44.589 deg) 3.55e-04 (69.2585 deg)
C1:LSC-REFL55_Q_ERR_DQ 1.84e-03 (16.3186 deg) 3.35e-03 (125.67 deg) 4.59e-05 (4.18718 deg)
C1:LSC-POX11_I_ERR_DQ 1.54e-01 (-157.852 deg) 6.07e-01 (-42.1078 deg) 5.55e-05 (73.3963 deg)
C1:LSC-POX11_Q_ERR_DQ 6.83e-05 (-148.591 deg) 6.37e-04 (121.983 deg) 1.35e-06 (43.7201 deg)
C1:LSC-POY11_I_ERR_DQ 1.85e-01 (36.1624 deg) 5.73e-01 (-43.1776 deg) 2.12e-04 (82.16 deg)
C1:LSC-POY11_Q_ERR_DQ 2.16e-05 (130.937 deg) 6.38e-05 (-173.194 deg) 1.40e-06 (47.5416 deg)

FPMI locked periods:
  - 1344129143 - 1344129520
  - 1344131106 - 1344131305
  - 1344133503 - 1344134020

Next:
- Restore CM servo for CARM

ELOG V3.1.3-